ย Consider any non-zero point in ๐‘…2 and name it (๐‘Ž,๐‘). Thenย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย 

(i). Write any four different paths that passes through your chosen point (๐‘Ž,๐‘).

(ii). Compute the limits of the following function when (๐‘ฅ,๐‘ฆ) โ†’ (๐‘Ž,๐‘) along all these four paths,ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย ย 

ย ๐‘“(๐‘ฅ,๐‘ฆ) = {

(๐‘ฅโˆ’๐‘Ž)2(๐‘ฆโˆ’๐‘) (๐‘ฅโˆ’๐‘Ž)4+(๐‘ฆโˆ’๐‘)2

, (๐‘ฅ,๐‘ฆ) โ‰  (๐‘Ž,๐‘) 0, (๐‘ฅ,๐‘ฆ) = (๐‘Ž,๐‘)ย ย 

(iii). Conclude from the results obtained in (ii) and answer whetherย lim (๐‘ฅ,๐‘ฆ)โ†’(๐‘Ž,๐‘) ๐‘“(๐‘ฅ,๐‘ฆ)ย exists or not.ย ย 

(iv). Is the function ๐‘“(๐‘ฅ,๐‘ฆ) continuous at the origin? Explain.ย ย 

(v). Also calculate ๐‘“๐‘ฅ(๐‘Ž,๐‘)ย and ๐‘“๐‘ฆ(๐‘Ž,๐‘).ย ย 

(vi). Write all points of differentiability of ๐‘“.


1
Expert's answer
2021-07-19T05:52:56-0400

i.

We consider "(a,b) = (0,0)"

Now the four different paths through "(0,0)" are

"y=x\\
y=x^2\\
y=x^3\\
y=x^4\\"


ii

"f(x,y) = \begin{cases}
 \frac{(x-a)^2(y-b)}{(x-a)^4(y-)^2} &\text{if } (x,y)\not=(a,b)\\
 0 &\text{if } (x,y)=(a,b)
\end{cases}\\\\
But (a,b) = (0,0)\\
f(x,y) = \begin{cases}
 \frac{x^2y}{x^4y^2} &\text{if } (x,y)\not=(0,0)\\
 0 &\text{if } (x,y)=(0,0)
\end{cases}\\\\
Now, \lim\nolimits_{(x,y) \to (0,0)}f(x,y)= \lim\nolimits_{(x,y) \to (0,0)}\frac{x^2y}{x^4y^2}\\
y=x\\
\lim\nolimits_{(x) \to (0)}\frac{x^2*x}{x^4*x^2}=0\\
y=x^2\\
\lim\nolimits_{(x) \to (0)}\frac{x^2*x^2}{x^4*x^4}=0.5\\
y=x^3\\
\lim\nolimits_{(x) \to (0)}\frac{x^2*x^3}{x^4*x^6}=0\\
y=x^4\\
\lim\nolimits_{(x) \to (0)}\frac{x^2*x^4}{x^4*x^8}=0\\"


iii.

We will conclude that along every path chosen the limit of the function exist.


iv.

The function is continuous at the origin along every path chosen except along "y=x^2" 7


v.

"f(x,y)=\frac{x^2y}{x^4+y^2}\\
f_n= \frac{∂f}{∂x}=\frac{(x^2+y^2)*2xy-x^2y(4x^3)}{(x^4+y^2)^2}\\
f_n(0,0)= Nonexistane\\
f_n= \frac{∂f}{∂y}=\frac{(x^4+y^2)*x^2-x^2y(2y)}{(x^4+y^2)^2}\\
f_n(0,0)= Nonexistane\\"

vi.

The function is differentiable at every point except at (0,0)